Esperantologio / Esperanto Studies 1 (1999), 5–6

Editorial Statement

Christer Kiselman

D ear reader:

Welcome to the first issue of a new journal! It will publish works on all aspects of the linguistic, social, and psychological phenomenon Esperanto. Its name is bilingual, to indicate that it will contain articles in Esperanto as well as in other languages. It appears in a print version as well as in an electronic version. In this way I hope to achieve maximum availability.

The journal is a kind of continuation: Between 1949 and 1961, Paul Neergaard published a journal with the name Esperantologio. A total of six issues were published.^{1} For a long time I was thinking about various ways to revive the journal. To guarantee high quality, I felt it would be essential to attract outstanding esperantologists and linguists to the idea. My proposal was welcomed by all the persons I contacted. They now form the Editorial Board, and I am most grateful to all of them for their support and for their efforts in the preparation of this first issue.

On the term esperantologio

The term esperantologio was introduced by Eugen Wüster in an article which appeared in 1921 in the journal Esperanto Triumfonta (later to be renamed Heroldo de Esperanto). In an article published in the journal Esperantologio he recounts the history of this term as well as that of interlingvistiko (Wüster 1955). For Wüster esperantologio denotes that branch of sinteza lingvistiko ‘synthetic linguistics’ which is applied to the already much used system of Esperanto. By sinteza lingvistiko he understands that relatively new branch of linguistics which not only observes but consciously influences (guides) the development of a language—nowadays we would say language planning. But he quotes other opinions as well, among them that of W. J. A. Manders, according to whom esperantologio is a science which observes and describes (Wüster 1955:211). The name of the present journal is to be understood in this sense.

The article La Esperantologio kaj ties disciplinoj, which Neergaard published in 1942, was important for early works on Esperanto. There he described in detail the various subbranches of Esperanto studies, and sketched the steps to be taken.

The policy of this journal

The phenomenon Esperanto is difficult to pinpoint. In my experience, it signifies very different things to different people. Many dreams are attached to it. The forces which inspire persons to learn and explore the language are of many kinds and hard to describe. Indeed it would be difficult to formulate a description of these forces

1The tables of contents of all issues is available on the web at the URL

http://www.math.uu.se/esperanto

on which people could agree. Maybe the most original and deepest of them is what Zamenhof pronounced in 1905: “[...] mi en la nuna momento ne estas ia naciano, sed simpla homo [...]” (at this moment I am not a member of any nation, but a simple human being; Zamenhof 1929:365).

This journal is based on scientific practice. Therefore we shall keep to Esperanto studies as an observing and describing science in the sense of Manders. The journal will not publish proposals to reform the language.

The policy of the journal is to study the existing linguistic and social phenomenon of Esperanto. However, this does not exclude the possibility that we might discuss in a coming issue what is or could be the objective of esperantologio.

On this issue

Since this journal is a kind of continuation of that edited by Neergaard, we devote some space to his life and work.

I am particularly happy that Marjorie Boulton permitted the inclusion of her paper in memory of Paul Neergaard. It is an essay prepared not only with a detailed knowledge of his personality but also with deep feelings concerning his whole life and work.

Carl Støp-Bowitz wrote soon after the death of Paul Neergaard a short essay on Neergaard’s work on scientific terminology.

The papers by Boulton and Støp-Bowitz were originally written for an attempt to revive Neergaard’s journal in 1987–1988. Humphrey Tonkin was kind enough to send me their manuscripts.

From Birthe Lapenna I received an article by Paul Neergaard himself, written in honor of Ivo Lapenna. She kindly consented to have it included here. In this way we can to some extent become acquainted with Neergaard’s thoughts on lexicology.

However, the three contributions just mentioned constitute a kind of exception due to their historical character. The four others are related to the linguistic and social aspects of the language.

Geraldo Mattos, the president of the Esperanto Academy, contributes an article on the Fundamento of Esperanto. This document from 1905 plays an important role in the language, in some measure replacing that of the long tradition of other languages.

Detlev Blanke describes the activities of an association of which he is the chairman.

Sabine Fiedler presents a study on phraseology, which is based on her extensive dissertation.

Finally we publish a pioneering study by Marc van Oostendorp on the phonology of Esperanto, more precisely on its syllable structure.

References

Paul Neergaard (1907–1987). 1942. La Esperantologio kaj ties disciplinoj. Taskoj kaj

rezultoj. In: Tra densa mallumo, pp. 37–64. Copenhagen: Laborista Esperanto Klubo. Eugen Wüster (1898–1977). 1955. La terminoj “esperantologio” kaj “interlingvistiko”. Es-

perantologio 1, No. 4, 209–214. L. L. Zamenhof (1859–1917). 1929. Originala verkaro. Antaŭparoloj – Gazetartikoloj –

Traktaĵoj – Paroladoj – Leteroj – Poemoj. Leipzig: Ferdinand Hirt & Sohn.